TABLE 3.3 Pro Forma Financial Statements for R&E Supplies, Inc., December 31, 2018 ($ thousands)

Income Statement

2018

Comments

Net sales

$25,766

25% increase

Cost of goods sold

  22,159

86% of sales

Gross profit

3,607

Expenses:

 General, selling, and administrative expenses

3,092

12% of sales

 Net interest expense

        90

Initially constant

Earnings before tax

425

Tax

      191

45% tax rate

Earnings after tax

$    234

Balance Sheet

Assets

Current assets:

 Cash and securities

$  1,271

18 days sales

 Accounts receivable

3,600

51-day collection period

 Inventories

2,462

9 times turnover

 Prepaid expenses

        20

Rough estimate

  Total current assets

7,353

Net fixed assets

      280

See text discussion

Total assets

$  7,633

Liabilities and Owners’ Equity

Current liabilities:

 Bank loan

$        0

 Accounts payable

3,582

59-day payables period

 Current portion of long-term debt

100

See text discussion

 Accrued wages

        22

Rough estimate

  Total current liabilities

3,704

Long-term debt

660

Common stock

150

Retained earnings

    1,697

See text discussion

Total liabilities and owners’ equity

$  6,211

External funding required

$  1,422

Table 3.3 shows the December 31, 2018 pro forma balance sheet and income statements for R&E Supplies, Inc. The pro-forma balance sheet shows that R&E Supplies will need external funding from the bank of $1.4 million. However, they show almost $1.3 million in cash and short-term securities. Why are they talking to the bank for such a large amount when they have most of this sum in their cash account?
Answer #1
The firm is probably trying to be too conservative so that they
can maintain a comfortable cash position.
Some of the ratios of liquidity that are relevant are:
Current ratio = 7353/3704 = 1.99
Quick ratio = (1271+3600)/3704 = 1.32
The above ratios indicate a comfortable liquidity position.
The cash cycle of the firm is reasonable at 51+365/9-59 = 33 days
But, the Days sales outstanding is high at 51, as also the days
payables outstanding of 59 days.
The firm can consider lowering the DSO and DSI and use a
part of the cash & marketable sexurities for financing the
fixed capital and working capital needs.
From the solvency angle, the firm's debt/equity ratio before taking the additional funding = (3704+660)/(150+1697) = 236%
This is very high and with the additional funding by bank, the
ratio will become (3704+660+1422)/(150+1697) = 313%
The bank may insist of enhancing the equity component of the
additional funding.